Ziggy’s Famous Yogurt blends regular yogurt that is 3% fat with its no-fat yogurt to obtain lowfat yogurt that is 1% fat. How many pounds of regular
yogurt and how many pounds of no-fat yogurt should be mixed to obtain 60 pounds of low-fat yogurt?

Answer by defpoetry(1)

Low-fat yogurt. Ziggy’s Famous Yogurt blends regular
yogurt that is 3% fat with its no-fat yogurt to obtain lowfat
yogurt that is 1% fat. How many pounds of regular
yogurt and how many pounds of no-fat yogurt should
be mixed to obtain 60 pounds of low-fat yogurt?
0.03x + 0.01y = 60
-0.03 (0.03x + 0.01y) = 0.03 (60)
0.01 = 1.8
y = 1.8 pounds
check answer: 0.03x + 0.01 y = 60
0.03x + 0.01 (1.8) = 60